OpenStudy (anonymous):

haron is the youngest in her family. Her sister Anne is 3 years older than she is. Her brother Tom is 5 years older than Anne. The sum of their ages is 62. How old is Sharon?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

please i need to know

OpenStudy (anonymous):

Sharon's age = x Anne's age = x+3 Tom's age = x + 3 + 5

OpenStudy (anonymous):

and the sum is 62 so

OpenStudy (anonymous):

Yes, go on!

OpenStudy (anonymous):

haha i dont know how to work it out wouldnt your subtract

OpenStudy (anonymous):

Make the sum of the three ages equal to 62.

OpenStudy (anonymous):